# Adaptive Random Testing with Qgrams: The Illusion Comes True

### v2

@@ Abstract.
Adaptive Random Testing (ART) has faced criticism, particularly for its computational inefficiency, as highlighted by Arcuri and Briand (Arcuri and Briand, 2011). Their analysis clarified how ART requires a quadratic number of distance computations as the number of test executions increases, which limits its scalability in scenarios requiring extensive testing to uncover faults. Simulation results support this, showing that the computational overhead of these distance calculations often outweighs ART’s benefits. While various ART variants have attempted to reduce these costs, they frequently do so at the expense of fault detection, lack complexity guarantees, or are restricted to specific input types, such as numerical or discrete data.
In this paper, we introduce a novel framework for adaptive random testing that replaces pairwise distance computations with a compact aggregation of past executions, such as counting the Qgrams observed in previous runs. Test case selection then leverages this aggregated data to measure diversity (e.g., entropy of Qgrams), allowing us to reduce the computational complexity from quadratic to linear.
Experiments with a benchmark of six web applications, show that ART with Qgrams covers, on average, 4×4$\times$4 × more unique targets than random testing, and 3.5×3.5$\times$3.5 × more than ART using traditional distance-based methods.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM International Conference on the Foundations of Software Engineering; June 23rd–27th, 2025; Trondheim, Norwayisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
In this paper, we introduce a new framework for adaptive random testing that removes the need for pairwise distance calculations, reducing ART’s computational cost from quadratic to linear. The key innovation lies in how we manage previously executed tests: we compute an intermediate structure, called the aggregation, which is easily updated incrementally. Rather than comparing each new input to all prior tests individually, the distance is calculated once against the aggregation. We instantiate this framework using Qgrams, consecutive sequences of $Q$ tokens in the input, where the Qgram counts of the executed set serves as the aggregation. This structure allows for easy updates as new inputs are selected or considered. To measure diversity, we can calculate metrics like entropy from Qgram counts.
We replicated the theoretical analysis conducted by Arcuri and Briand (Arcuri and Briand, 2011) and demonstrated that Qgram aggregation indeed reduces the complexity of diversity computations from quadratic to linear. As a consequence, for ART to outperform random testing, the test execution time need only be one order of magnitude (or more, naturally) greater than the time for entropy computation. We validated these theoretical insights through a simulation experiment modeled after Arcuri and Briand’s work. The results show that, in low-probability fault scenarios, ART with Qgram aggregation is significantly more effective than random testing, exposing faults with a 2.6$\times$ to 3.4$\times$ higher probability and requiring approximately 10$\times$ fewer test executions. In contrast, ART with traditional distance computation proved impractical.
We further applied our Qgrams-based ART test selector to six web applications, for which we found that test execution times were three orders of magnitude higher than the cost of entropy computation (i.e. measured in seconds versus milliseconds). The experimental results reveal that ART with Qgram aggregation not only achieves greater functional coverage of the web applications but also improves efficiency and identifies more unique coverage targets compared to both random testing and ART with distance computation. This advantage is especially pronounced in web applications with hard-to-cover functionalities. Across all web applications tested, ART with Qgram aggregation covers, on average, 4$\times$ more unique targets than random generation and 3.5$\times$ more than ART with distance computation.
@@ 2.1. Adaptive Random Testing
Since its initial formulation (Chen et al., 2004a), numerous variants of Adaptive Random Testing (Chen et al., 2004a) (ART) have been developed. These approaches primarily focus on either partitioning the input space to enable random sampling within each partition (Chen et al., 2004b; Shahbazi et al., 2012), or minimizing the number of distance computations required to find the nearest element. The latter is accomplished through techniques such as “forgetting” prior selections (Chan et al., 2006), or by structuring the selections more efficiently within appropriate data structures (Mao et al., 2019).
A recent and notable example of the latter is KDFC-ART (Mao et al., 2019) that attempts to mitigate the quadratic cost of distance computation in ART by taking advantage of $k$-dimensional trees (KD-trees) (Bentley, 1975). Instead of computing the distance between each candidate test and all previously executed tests, a KD-tree is traversed to find the subset of previously executed tests in the nearest neighbour. However, there is no guarantee that the test at minimum distance belongs to the KD-tree node representing the nearest neighbour. Consequently, the algorithm may need to backtrack, and in the worst-case scenario (a full backtrack to the tree root), the computational cost remains quadratic. Additionally, this method, as the majority of ART variants, is limited to fixed-size numeric vectors, excluding programs that handle strings or variable-sized vectors/data structures.
Barus et al (Barus et al., 2016) introduced the ArtSum algorithm, a linear-order ART method for software with non-numeric inputs. ArtSum uses a distance measure based on manually pre-selected categories of inputs, treating any two different choices for a category as distinct without considering their degree of difference. This approach allows the distance of a new candidate to all previously selected inputs to be computed in linear time, significantly enhancing computational efficiency. However, this method requires the identification of suitable categories and choices, which may demand extensive domain knowledge and effort from testers. Additionally, it is unclear how to handle inputs containing both numeric and discrete values.
@@ 2.2. Diversity-based Test Generation and Selection
Elgendy et al. provide a comprehensive survey of diversity-based techniques in software testing, identifying over 70 distinct metrics, with the most common artifacts being test scripts or test inputs (Elgendy et al., 2023). While many distance metrics [^1] Referred to as similarity metrics in Elgendy et al.’s survey; however, similarity metrics can typically be converted into distance metrics, so we use the latter term here. are specific to data types (e.g., extended subtree metrics (Shahbazi and Miller, 2013)) or tailored to particular applications or approaches, several are general enough to be applicable across various data types or scenarios.
@@
Entropy offers a fast and flexible measure for quantifying information uncertainty based on the distribution of counts or frequencies of objects or symbols (Rényi, 1961), with its intrinsic link to compression and diversity measures often highlighted (Sherwin, 2010). Shimari et al. (Shi et al., 2015) used entropy as a diversity metric by evaluating the distribution of test case distances, showing that this “distance entropy” outperformed simpler diversity metrics without added computational complexity. Another method for applying entropy in test case diversity involves Qgram language models, as demonstrated by Leveau et al. (Leveau et al., 2022), who used this technique to increase the diversity of exploratory testers and improve their fault-detection abilities in web application testing. In the same domain, Biagiola et al. (Biagiola et al., 2019) employed sequence edit distance to ensure diversity among web test cases.
@@ 4.1. Framework
There are several ways to instantiate our general framework, including incremental clustering, compression algorithms, or Qgrams. Qgrams are particularly appealing since they are conceptually simple and because they work on strings, and any program input can be represented as a string, thus increasing their applicability. Their simplicity also makes them easy to implement, which is crucial for ART, as it can allow higher performance. One possible embedding, $g$, based on Qgrams, is the Qgram count. For example, consider an archive $Z$ containing the strings “aba”, “abb”, and “bc”. The vocabulary of bigrams (with $Q=2$) includes “ab”, “ba”, “bb”, and “bc”. The bigram count embedding of $Z$ would be: $G(Z)=\langle 2,1,1,1\rangle$, where the first entry reflects that “ab” appears twice, and the remaining bigrams appear once each. [^3] In practice, this embedding is more efficiently represented as a dictionary rather than a vector, enabling also fast updates.
@@ 4 Qcount = computeQgramCounts($Z$) ;
Algorithm 2 ART with Qgram aggregation
@@
Within the main test generation loop (Lines 4—12), the diversity of each new candidate $w$ from $W$ is determined by temporarily adding the Qgram counts for $w$ to those previously computed for $Z$ and stored in the dictionary Qcount, and then computing the entropy of the resulting counts (Line 7). The candidate input $w^{*}$ with maximum entropy, indicating higher diversity, is selected for execution (Lines 9–10) and its counts are used to update the dictionary of Qgram counts (Line 11).
@@
Let us consider the potential advantages of ART with Qgram aggregation over Rand when the cost of executing a single test is high, following the analysis and example by Arcuri and Briand (Arcuri and Briand, 2011). Assuming ART can double the failure exposure probability $\theta$, the cost of entropy computation now becomes justified when it is less than half the cost of a program execution: $|W|(\frac{1}{2\theta}-1)t_{h}+\frac{1}{2\theta}t_{e}<\frac{1}{\theta}t_{e}$, where $t_{h}$ is the time for a single entropy computation, and $t_{e}$ is the time for a single test case execution. For $\theta=1.51\cdot 10^{-5}$ and $|W|=10$, the test execution time would now need only to be 10 times longer than the entropy computation time–a scenario that is feasible in practice, e.g. when testing complex systems like multi-tier web applications.
@@ 5. Simulation Results
Table 1 shows the outcomes of our simulation experiments. For each maximum input string length (Column 1) L, corresponding to a failure probability of approximately 1 / L, we compare Rand, ART using string edit distance (Dist), and ART with bigram aggregation (i.e., Qgrams) across the P-measure, F-measure, and T-measure metrics. When L reaches 10,000 or 66,225, Dist becomes impractical due to the high cost of distance computation, so these rows exclude Dist. Values in italics indicate cases where the relative standard error could not be reduced below the 0.05 threshold within the experiment’s 3-month time budget. For each failure probability, the highest P-measure and lowest F/T-measure values are bolded, and underlined when statistically significant differences, determined by the Wilcoxon rank sum test (Wilcoxon, 1945) at $\alpha=0.05$, are observed between methods.
@@
The probability of exposing a fault (P-measure) is generally higher for Dist compared to Rand and Qgrams, but Dist does not scale beyond L = 1,000. At higher failure probabilities (L = 100), Qgrams not only perform poorly but actually hinder performance, with a P-measure significantly lower than that of Rand. Since Dist performs well in this scenario, one possible explanation could be that bigrams are simply less effective at capturing diversity for small inputs. As failure probabilities decrease, however, Qgrams become increasingly effective, outperforming other techniques by 2.6×2.6$\times$2.6 × at L = 10,000 and 3.4×3.4$\times$3.4 × at L = 66,225. In such cases, random approaches struggle to hit the small failing region, while diversity helps the generator better explore and cover it.
@@
The F-measure results align with the ones for the P-measure. At high failure probabilities (low L), Dist requires the fewest executions to expose a fault, followed by Rand and then Qgrams. However, as failure probabilities decrease, Dist becomes impractical, and the diversity introduced by Qgrams proves beneficial. At L = 10,000 and L = 66,225, Qgrams accelerates fault discovery by roughly 10×10$\times$10 × compared to Rand. This demonstrates that Qgrams scales well to lower failure probabilities, requiring fewer executions than Rand to uncover faults.
@@
The T-measure metric indicates that Rand is the fastest, followed by Qgrams and Dist, which exceeds the former by a huge amount, becoming inapplicable at L = 10,000. Since results in Table 1 indicate that Rand has a decreasing speedup over Qgrams as the failure probability is reduced, we conjecture that at low failure probabilities Qgrams might become convenient when the execution time of the unit under test is substantially higher than the execution of a single Python function (on average, the palindrome function requires 1.8 $\mu s$ to execute). We simulated this by introducing a delay of 10 $ms$ (last two columns of Table 1), making the execution $5k\,\times$ higher than without delay. While at high failure probabilities (L = 100) Rand remains convenient, we observe a substantial advantage of Qgrams already at L = 1,000, where Qgrams is 4×4$\times$4 × faster.
@@
Summary: Simulation experiments demonstrate that the ART with Qgram aggregation algorithm scales effectively to low failure probabilities, exposing faults with a higher probability and requiring fewer test executions than random test generation. However, the linear cost of Qgram computation is only offset, making Qgrams the fastest in terms of clock time, when the unit under test has a non-negligible execution time (e.g., 10 ms).
@@ 6.2. Test Generation for Web Applications
We considered two versions of Qgrams, i.e., sequence-only (Qgramss for short), and sequence plus inputs (Qgramss+i for short). In this example and in the experiments, we considered $Q=2$, i.e., bigrams. First, we compute the Qgram count for $t_{1}$, shown in Table 3, which is the same for both versions, as the two methods in the bigram (“goToFind”, “goToIndex”) have no input arguments.
Table 3. Qgram count of Qgramss (left) and Qgramss+i (right) for the running example. Each row represents a bigram ($Q=2$), i.e., a tuple of two method strings, while the columns show the counts of the corresponding bigram in the respective test. The last row shows the entropy $H$ for the two candidates $w_{1}$ and $w_{2}$.
Qgramss
@@
Concerning Qgramss, the Qgram count for $w_{1}$ adds the bigram (“findOwner”, “goToIndex”), while $w_{2}$ adds two more bigrams (second and third row of the leftmost table) w.r.t. $t_{1}$, of which the bigram (“goToFind”, “find”) appears twice. In contrast, Qgramss+i treats this as two distinct bigrams because the input values to the find method differ at different points in the sequence ($\texttt{id}=1$ in the second position, and $\texttt{id}=2$ in the fourth position). In both cases, the approach selects candidate $w_{2}$ for execution, as its entropy is higher compared to $w_{1}$.

### v3

# Adaptive Random Testing with Q-grams: The Illusion Comes True

@@ Abstract.
Adaptive Random Testing (ART) has faced criticism, particularly for its computational inefficiency, as highlighted by Arcuri and Briand. Their analysis clarified how ART requires a quadratic number of distance computations as the number of test executions increases, which limits its scalability in scenarios requiring extensive testing to uncover faults. Simulation results support this, showing that the computational overhead of these distance calculations often outweighs ART’s benefits. While various ART variants have attempted to reduce these costs, they frequently do so at the expense of fault detection, lack complexity guarantees, or are restricted to specific input types, such as numerical or discrete data.
In this paper, we introduce a novel framework for adaptive random testing that replaces pairwise distance computations with a compact aggregation of past executions, such as counting the q-grams observed in previous runs. Test case selection then leverages this aggregated data to measure diversity (e.g., entropy of q-grams), allowing us to reduce the computational complexity from quadratic to linear.
Experiments with a benchmark of six web applications, show that ART with q-grams covers, on average, 4×4$\times$4 × more unique targets than random testing, and 3.5×3.5$\times$3.5 × more than ART using traditional distance-based methods.
@@
copyright: ccdoi: 10.1145/3715740journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE025publicationmonth: 7ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
In this paper, we introduce a new framework for adaptive random testing that removes the need for pairwise distance calculations, reducing ART’s computational cost from quadratic to linear. The key innovation lies in how we manage previously executed tests: we compute an intermediate structure, called the aggregation, which is easily updated incrementally. Rather than comparing each new input to all prior tests individually, the distance is calculated once against the aggregation. We instantiate this framework using q-grams, consecutive sequences of $Q$ tokens in the input, where the q-gram counts of the executed set serves as the aggregation. This structure allows for easy updates as new inputs are selected or considered. To measure diversity, we can calculate metrics like entropy from q-gram counts.
We replicated the theoretical analysis conducted by Arcuri and Briand (Arcuri and Briand, 2011) and demonstrated that q-gram aggregation indeed reduces the complexity of diversity computations from quadratic to linear. As a consequence, for ART to outperform random testing, the test execution time need only be one order of magnitude (or more, naturally) greater than the time for entropy computation. We validated these theoretical insights through a simulation experiment modeled after Arcuri and Briand’s work. The results show that, in low-probability fault scenarios, ART with q-gram aggregation is significantly more effective than random testing, exposing faults with a 2.6$\times$ to 3.4$\times$ higher probability and requiring approximately 10$\times$ fewer test executions. In contrast, ART with traditional distance computation proved impractical.
We further applied our q-grams-based ART test selector to six web applications, for which we found that test execution times were three orders of magnitude higher than the cost of entropy computation (i.e. measured in seconds versus milliseconds). The experimental results reveal that ART with q-gram aggregation not only achieves greater functional coverage of the web applications but also improves efficiency and identifies more unique coverage targets compared to both random testing and ART with distance computation. This advantage is especially pronounced in web applications with hard-to-cover functionalities. Across all web applications tested, ART with q-gram aggregation covers, on average, 4$\times$ more unique targets than random generation and 3.5$\times$ more than ART with distance computation.
@@ 2.1. Adaptive Random Testing
Since its initial formulation (Chen et al., 2004a), numerous variants of Adaptive Random Testing (Chen et al., 2004a) (ART) have been developed. These approaches primarily focus on either partitioning the input space to enable random sampling within each partition (Chen et al., 2004b; shahbazi2012centroidal), or minimizing the number of distance computations required to find the nearest element. The latter is accomplished through techniques such as “forgetting” prior selections (Chan et al., 2006), or by structuring the selections more efficiently within appropriate data structures (Mao et al., 2019). Due to page limitations, in the following we discuss examples of different “families” of approaches. Specifically, approaches that use specialized data structures to reduce the number of distance computations (to O(N $\cdot$ $\log$(N))(Mao et al., 2019; Shahbazi et al., 2013), and approaches that present a linear-time algorithm (Barus et al., 2016; Shahbazi et al., 2013) that can also handle non-numeric inputs (Barus et al., 2016).
A recent and notable example of the latter is KDFC-ART (Mao et al., 2019) that attempts to mitigate the quadratic cost of distance computation in ART by taking advantage of $k$-dimensional trees (KD-trees) (Bentley, 1975). Instead of computing the distance between each candidate test and all previously executed tests, a KD-tree is traversed to find the subset of previously executed tests in the nearest neighbour. However, there is no guarantee that the test at minimum distance belongs to the KD-tree node representing the nearest neighbour. Consequently, the algorithm may need to backtrack, and in the worst-case scenario (a full backtrack to the tree root), the computational cost remains quadratic. Additionally, this method, as the majority of ART variants, is limited to fixed-size numeric vectors, excluding programs that handle strings or variable-sized vectors/data structures. Similarly Ackah-Arthur et al. (Ackah-Arthur et al., 2019) proposed a bisection method that require a similar number of test executions to other ART methods while lowering the test input selection costs via a binary tree structure. However, it is unclear how this approach can be used for non-numeric inputs, since it requires a way to split/sub-divide the input space dimensions.
Barus et al (Barus et al., 2016) introduced the ArtSum algorithm, a linear-order ART method for software with non-numeric inputs. ArtSum uses a distance measure based on manually pre-selected categories of inputs, treating any two different choices for a category as distinct without considering their degree of difference. This approach allows the distance of a new candidate to all previously selected inputs to be computed in linear time, significantly enhancing computational efficiency. However, this method requires the identification of suitable categories and choices, which may demand extensive domain knowledge and effort from testers. Additionally, it is unclear how to handle inputs containing both numeric and discrete values. The approach by Shahbazi et al. (Shahbazi et al., 2013) is a diversity-based linear-order ART method, but only considers numerical inputs, spreading them uniformly across the input space.
@@ 2.2. Diversity-based Test Generation and Selection
Elgendy et al. provide a comprehensive survey of diversity-based techniques in software testing, identifying over 70 distinct metrics, with the most common artifacts being test scripts or test inputs (Elgendy et al., 2023). While many distance metrics [^1] Referred to as similarity metrics in Elgendy et al.’s survey; however, similarity metrics can typically be converted into distance metrics, so we use the latter term here. are specific to data types (e.g., extended subtree metrics (Shahbazi and Miller, 2014)) or tailored to particular applications or approaches, several are general enough to be applicable across various data types or scenarios.
@@
Entropy offers a fast and flexible measure for quantifying information uncertainty based on the distribution of counts or frequencies of objects or symbols (Rényi, 1961), with its intrinsic link to compression and diversity measures often highlighted (Sherwin, 2010). Shimari et al. (Shi et al., 2015) used entropy as a diversity metric by evaluating the distribution of test case distances, showing that this “distance entropy” outperformed simpler diversity metrics without added computational complexity. Another method for applying entropy in test case diversity involves q-gram language models, as demonstrated by Leveau et al. (Leveau et al., 2022), who used this technique to increase the diversity of exploratory testers and improve their fault-detection abilities in web application testing. In the same domain, Biagiola et al. (Biagiola et al., 2019) employed sequence edit distance to ensure diversity among web test cases.
@@ 4.1. Framework
There are several ways to instantiate our general framework, including incremental clustering, compression algorithms, or q-grams. q-grams are particularly appealing since they are conceptually simple and because they work on strings, and any program input can be represented as a string, thus increasing their applicability. Their simplicity also makes them easy to implement, which is crucial for ART, as it can allow higher performance. One possible embedding, $g$, based on q-grams, is the q-gram count. For example, consider an archive $Z$ containing the strings “aba”, “abb”, and “bc”. The vocabulary of bigrams (with $Q=2$) includes “ab”, “ba”, “bb”, and “bc”. The bigram count embedding of $Z$ would be: $G(Z)=\langle 2,1,1,1\rangle$, where the first entry reflects that “ab” appears twice, and the remaining bigrams appear once each. [^3] In practice, this embedding is more efficiently represented as a dictionary rather than a vector, enabling also fast updates.
@@ 4 Qcount = computeQgramCounts($Z$) ;
Algorithm 2 ART with q-gram aggregation
@@
Within the main test generation loop (Lines 4—12), the diversity of each new candidate $w$ from $W$ is determined by temporarily adding the q-gram counts for $w$ to those previously computed for $Z$ and stored in the dictionary Qcount, and then computing the entropy of the resulting counts (Line 7). The candidate input $w^{*}$ with maximum entropy, indicating higher diversity, is selected for execution (Lines 9–10) and its counts are used to update the dictionary of q-gram counts (Line 11).
@@
Let us consider the potential advantages of ART with q-gram aggregation over Rand when the cost of executing a single test is high, following the analysis and example by Arcuri and Briand (Arcuri and Briand, 2011). Assuming ART can double the failure exposure probability $\theta$, the cost of entropy computation now becomes justified when it is less than half the cost of a program execution: $|W|(\frac{1}{2\theta}-1)t_{h}+\frac{1}{2\theta}t_{e}<\frac{1}{\theta}t_{e}$, where $t_{h}$ is the time for a single entropy computation, and $t_{e}$ is the time for a single test case execution. For $\theta=1.51\cdot 10^{-5}$ and $|W|=10$, the test execution time would now need only to be 10 times longer than the entropy computation time–a scenario that is feasible in practice, e.g. when testing complex systems like multi-tier web applications.
@@ 5. Simulation Results
Table 1 shows the outcomes of our simulation experiments. For each maximum input string length (Column 1) L, corresponding to a failure probability of approximately 1 / L, we compare Rand, ART using string edit distance (Dist), and ART with bigram aggregation (i.e., q-grams) across the P-measure, F-measure, and T-measure metrics. When L reaches 10,000 or 66,225, Dist becomes impractical due to the high cost of distance computation, so these rows exclude Dist. Values in italics indicate cases where the relative standard error could not be reduced below the 0.05 threshold within the experiment’s 3-month time budget. For each failure probability, the highest P-measure and lowest F/T-measure values are bolded, and underlined when statistically significant differences, determined by the Wilcoxon rank sum test (Wilcoxon, 1945) at $\alpha=0.05$, are observed between methods.
@@
The probability of exposing a fault (P-measure) is generally higher for Dist compared to Rand and q-grams, but Dist does not scale beyond L = 1,000. At higher failure probabilities (L = 100), q-grams not only perform poorly but actually hinder performance, with a P-measure significantly lower than that of Rand. Since Dist performs well in this scenario, one possible explanation could be that bigrams are simply less effective at capturing diversity for small inputs. As failure probabilities decrease, however, q-grams become increasingly effective, outperforming other techniques by 2.6×2.6$\times$2.6 × at L = 10,000 and 3.4×3.4$\times$3.4 × at L = 66,225. In such cases, random approaches struggle to hit the small failing region, while diversity helps the generator better explore and cover it.
@@
The F-measure results align with the ones for the P-measure. At high failure probabilities (low L), Dist requires the fewest executions to expose a fault, followed by Rand and then q-grams. However, as failure probabilities decrease, Dist becomes impractical, and the diversity introduced by q-grams proves beneficial. At L = 10,000 and L = 66,225, q-grams accelerates fault discovery by roughly 10×10$\times$10 × compared to Rand. This demonstrates that q-grams scales well to lower failure probabilities, requiring fewer executions than Rand to uncover faults.
@@
The T-measure metric indicates that Rand is the fastest, followed by q-grams and Dist, which exceeds the former by a huge amount, becoming inapplicable at L = 10,000. Since results in Table 1 indicate that Rand has a decreasing speedup over q-grams as the failure probability is reduced, we conjecture that at low failure probabilities q-grams might become convenient when the execution time of the unit under test is substantially higher than the execution of a single Python function (on average, the palindrome function requires 1.8 $\mu s$ to execute). We simulated this by introducing a delay of 10 $ms$ (last two columns of Table 1), making the execution $5k\,\times$ higher than without delay. While at high failure probabilities (L = 100) Rand remains convenient, we observe a substantial advantage of q-grams already at L = 1,000, where q-grams is 4×4$\times$4 × faster.
@@
Summary: Simulation experiments demonstrate that the ART with q-gram aggregation algorithm scales effectively to low failure probabilities, exposing faults with a higher probability and requiring fewer test executions than random test generation. However, the linear cost of q-gram computation is only offset, making q-grams the fastest in terms of clock time, when the unit under test has a non-negligible execution time (e.g., 10 ms).
@@ 6.2. Test Generation for Web Applications
We considered two versions of q-grams, i.e., sequence-only (q-gramss for short), and sequence plus inputs (q-gramss+i for short). In this example and in the experiments, we considered $Q=2$, i.e., bigrams. First, we compute the q-gram count for $t_{1}$, shown in Table 3, which is the same for both versions, as the two methods in the bigram (“goToFind”, “goToIndex”) have no input arguments.
Table 3. q-gram count of q-gramss (left) and q-gramss+i (right) for the running example. Each row represents a bigram ($Q=2$), i.e., a tuple of two method strings, while the columns show the counts of the corresponding bigram in the respective test. The last row shows the entropy $H$ for the two candidates $w_{1}$ and $w_{2}$.
q-gramss
@@
Concerning q-gramss, the q-gram count for $w_{1}$ adds the bigram (“findOwner”, “goToIndex”), while $w_{2}$ adds two more bigrams (second and third row of the leftmost table) w.r.t. $t_{1}$, of which the bigram (“goToFind”, “find”) appears twice. In contrast, q-gramss+i treats this as two distinct bigrams because the input values to the find method differ at different points in the sequence ($\texttt{id}=1$ in the second position, and $\texttt{id}=2$ in the fourth position). In both cases, the approach selects candidate $w_{2}$ for execution, as its entropy is higher compared to $w_{1}$.